OpenStudy (zubhanwc3):

Which of the following expressions does NOT have the same derivative as y=log_3 x

OpenStudy (zubhanwc3):

Which of the following expressions does not have the same derivative as \[y=\log_{3} x\] a)\[\frac{ 1 }{ 2 }\log_{6}x \] b)\[\log_{9}x^2, x >0\] My guess is B, but i dont see how a has the same derivative as \[y=\log_{3} x\]

OpenStudy (anonymous):

Do you know what the derivative of \(\log_3x\) is?

OpenStudy (zubhanwc3):

derivative of lnx/ln3

OpenStudy (zubhanwc3):

which should be 3/x i think

OpenStudy (anonymous):

Hmm, not quite. \[\frac{d}{dx}\log_3x=\frac{d}{dx}\frac{\ln x}{\ln 3}=\frac{1}{x\ln3}\]

OpenStudy (zubhanwc3):

o wait, i get it, b is 6/x then u divide that by 2, and u get 3/x?

OpenStudy (anonymous):

Your derivatives aren't right, but your reasoning is.

OpenStudy (zubhanwc3):

>.> forgot that ln3 is a constant multiple >.>

OpenStudy (zubhanwc3):

the derivative of a should be 1/2xln6 and derivative of b should be 2/xln9 correct?

OpenStudy (anonymous):

\[y=\frac{1}{2}\log_6x~~\implies~~y'=\frac{1}{2}\frac{\ln x}{\ln6}=\frac{\ln x}{\ln36}\] Hmm looks like (a) isn't right.

OpenStudy (zubhanwc3):

in my opinion, neither have the same derivative as log_3 of x how'd u get ln36 on the bottom btw?

OpenStudy (anonymous):

It's a property of logarithms. \(a\log_bx=\log_bx^a\).

OpenStudy (zubhanwc3):

but how does that get u ln 36?

OpenStudy (anonymous):

In the denominator, \[2\ln6=\ln6^2=\ln36\]

OpenStudy (zubhanwc3):

i get it, and just to confirm, neither of these have the same derivative as log_3 x?

OpenStudy (anonymous):

\[y=\log_9x^2=\frac{\ln x^2}{\ln9}=\frac{2\ln x}{\ln 9}~~\implies~~y'=\frac{2}{x\ln9}=\frac{1}{\frac{1}{2}x\ln 9}=\frac{1}{x\ln9^{1/2}}\] Nope, it's (b).

OpenStudy (zubhanwc3):

how do u get the 1/2 on the bottom?

OpenStudy (zubhanwc3):

by multiplying the top and bottom by 1/2?